OpenStudy (polymorphism):

How to simplify this ? I will post what I have in a minute. \[ \frac{\sqrt(6)}{6 + \sqrt(x)} \]

OpenStudy (polymorphism):

\[ \frac{\sqrt{x}}{6 + \sqrt{x}}\] \[ \frac{\sqrt{x}}{6 + \sqrt{x}} * \frac{{6 - \sqrt{x}}}{6 - \sqrt{x}} = \frac{{6 \sqrt{6}- \sqrt{6}\sqrt{x}}}{36 - x}\] Is that simplified?

OpenStudy (polymorphism):

Sorry, it i suppose to be \[ \frac{\sqrt{6}}{6 + \sqrt{x}} \] and then \[ \frac{\sqrt{6}}{6 + \sqrt{x}} * \frac{{6 - \sqrt{x}}}{6 - \sqrt{x}} = \frac{{6 \sqrt{6}- \sqrt{6}\sqrt{x}}}{36 - x} \]

OpenStudy (lannyxx):

looks good to me. the rule to simplifying these is as you've done, just make sure no roots exist in the denominator. it's a headache

OpenStudy (polymorphism):

Ty